# The Algorithmic Gaze: An Audit and Ethnography of the LAION-Aesthetics Predictor Model

### v2

@@ 5.2.1. Pluralistic Aesthetic Evaluation
As discussed above, we found numerous aesthetic in the behavior and design of LAP. In raising these issues, we do not wish advocate for designing “better” universal aesthetic quality evaluation models and datasets. As assessing the visual aesthetic quality of an image is deeply subjective and value-laden (Bourdieu, 1984), any attempt to represent the aesthetic quality of an image with a single number will be deeply flawed. As an alternative to this normative problem formulation, we encourage researchers to make use of more descriptive evaluating metrics (Röttger et al., 2022). Although LAP is often used as a general measure of aesthetic quality (Ding et al., 2024; Fong and See, 2024; Gao and Liu, 2024; He et al., 2023), we found that it could be more accurately described as a measure of photorealism. As described above, there are a number of limitations to using realism to assess the quality of an image. However, a model that measures photorealism is not in-and-of-itself harmful. Rather, we caution researchers against treating photorealism as a prescriptive rather than descriptive property of images, conflating how images ought to be with merely how images can be.
@@
To support more pluralistic AI alignment (Sorensen et al., 2024b), we encourage the use of descriptive measurements when evaluating visual generative AI model outputs and training data. Abandoning the flawed construct of universal “aesthetics” would allow researchers to more clearly articulate which aesthetic values their visual generative AI models prioritize, such as photorealism or particular visual cultures. For instance, Tencent’s Hunyuan-DiT T2I model is specifically designed to improve Chinese visual element understanding (Li et al., 2024b). In addition to improving transparency, more descriptive evaluations may help artists choose models that better align their tastes, perhaps addressing artists’ critique of increasing photorealistic T2I models (Shelby et al., 2024; Taylor et al., 2025). The seeds of more descriptive measures of aesthetic quality can actually be seen in the 2012 AVA dataset itself (Murray et al., 2012). Recognizing the limits of de-contextualized aesthetic evaluation, Murray et al. grouped similar photography challenges to measure 14 different photographic styles, such as “motion blur” and “long exposure.”
@@
The powerful companies developing T2I models may have little incentive to shift from prescriptive to descriptive aesthetic evaluations. For-profit companies go to great lengths arguing that their new models supplant their competitors. For instance, Meta researchers manipulated benchmarks to give an inflated impression of Llama 4’s performance (Cranz, 2025). Acknowledging that different T2I models may be better in different visual contexts undermines companies’ efforts to prove that their models are uniquely “state-of-the-art.” Much like corporate AI rhetoric, academic AI research is increasingly driven by benchmark competition (Gururaja et al., 2023; Orr and Kang, 2024), often without questioning the foundational assumptions embedded in said benchmarks (Raji et al., 2021). We encourage future researchers to examine what the universal aesthetics measurements used by companies and AI researchers actually measure.
5.2.2. Combining an Audit and Ethnography
@@
In this work, we were able to examine upstream design decisions contributing to downstream biases by combining an audit and ethnography. Auditing is a core FAccT methodology, with the regex “audit*” appearing in nearly half of all FAccT publications. [^12] https://dl.acm.org/action/doSearch?fillQuickSearch=false&target=advanced&ConceptID=122580&expand=all&AllField=audit* With some notable exceptions (Miceli et al., 2021; Marda and Narayan, 2020; Riley, 2024; Scheuerman, 2024; Passi and Barocas, 2019; Baack, 2024a), ethnography is far less common at FAccT. As Seaver (2017) argues, audits can help identify the existence of algorithmic biases while ethnographies can help explain where these biases may have originated by focusing on the people behind the algorithm. In addition to focusing on the harms of AI models themselves, responsible AI researchers are increasingly advocating for “studying up” the powerful people behind AI development and deployment (Barabas et al., 2020; Kawakami et al., 2024). In service of these goals, we advocate for combining audit and ethnography as a socio-technical method for evaluating algorithmic systems.
@@
While one does not need permission from its creators to audit an algorithm, ethnographic methods often require negotiating access (Scheuerman, 2024). For example, prior ethnographic FAccT research has relied on access to field sites like major tech companies (Scheuerman, 2024; Passi and Barocas, 2019) or the Delhi Police Headquarters (Marda and Narayan, 2020). However, maintaining access can limit critical scholarship (Scheuerman, 2024). Even in the case of “open” models, development practices may be opaque (Widder et al., 2024). We were only able to understand out how LAP was made by finding an explainer video in an old Discord message. Herein lies a lesson for future ethnographic FAccT research beyond traditional field sites. Ethnographies of algorithms may require “scavenging” (Seaver, 2017) for diverse materials, including formal documentation, datasets, academic papers, and social media conversations. At the same time, sharing an unlisted YouTube video—even if posted in a public Discord—could raise ethical concerns. We have chosen to include this video in our analysis due to LAION’s advocacy for data scraping and refusal to remove artists’ work from their datasets (Filikhina and Elsaß, 2025). That said, the “public-ness” of data does not always make its analysis ethical (Zimmer, 2010).
@@ 6. Conclusion
Generative AI models and algorithmic systems are increasingly mediating our media, art, and culture. Therefore, it is imperative to understand for whose taste these systems are designed. In this work, we audited the LAION-Aesthetics Predictor (LAP) model, finding that LAP reflects imperial, realist, and male aesthetic sensibilities. Through an ethnographic analysis of public materials, we found that the development of LAP reflects similar biases. In light of these findings, we call for greater attention on aesthetic biases as a source of representational harm and advocate for more pluralistic revelations. Much like we examine the algorithmic gaze of LAP, we encourage future research into the cultural consequences of algorithmic “quality” assessment.
@@ Acknowledgements.
Names removed for review

### v3

@@ 5.2.1. Pluralistic Aesthetic Evaluation
We found numerous aesthetic biases in the behavior and design of LAP. However, in raising these issues, we do not advocate for designing “better” universal aesthetic evaluation models or datasets. As assessing the visual aesthetic quality of an image is deeply subjective and value-laden (Bourdieu, 1984), any attempt to represent the aesthetic quality of an image with a single number will be deeply flawed. As an alternative to this normative problem formulation, we encourage researchers to make use of more descriptive evaluating metrics (Röttger et al., 2022). Although LAP is often used as a general measure of aesthetic quality (Ding et al., 2024; Fong and See, 2024; Gao and Liu, 2024; He et al., 2023), we found that it could be more accurately described as a measure of photorealism. As described above, there are a number of limitations to using realism to assess the quality of an image. However, a model that measures photorealism is not in-and-of-itself harmful. Rather, we caution researchers against treating photorealism as a prescriptive rather than descriptive property of images, conflating how images ought to be with merely how images can be.
@@
To support more pluralistic AI alignment (Sorensen et al., 2024b), we encourage the use of descriptive measurements when evaluating visual generative AI model outputs and training data. Abandoning the flawed construct of universal “aesthetics” would allow researchers to more clearly articulate which aesthetic values their visual generative AI models prioritize, such as photorealism or particular visual cultures. For instance, Tencent’s Hunyuan-DiT T2I model is specifically designed to improve Chinese visual element understanding (Li et al., 2024b). In addition to improving transparency, more descriptive evaluations may help artists choose models that better align with their tastes, perhaps addressing artists’ critique of increasing T2I photorealism (Shelby et al., 2024; Taylor et al., 2025). The seeds of more descriptive aesthetic evaluation can actually be seen in the 2012 AVA dataset itself (Murray et al., 2012). Recognizing the limits of de-contextualized aesthetic evaluation, Murray et al. grouped similar photography challenges to measure 14 different photographic styles, such as “motion blur” and “long exposure.”
@@
The companies developing T2I models may have little incentive to shift from prescriptive to descriptive aesthetic evaluations. For-profit companies go to great lengths arguing that their models supplant competitors’ models. For instance, Meta researchers manipulated benchmarks to give an inflated impression of Llama 4’s performance (Cranz, 2025). Acknowledging that different T2I models may be better in different contexts undermines companies’ efforts to prove that their models are uniquely “state-of-the-art.” Much like corporate AI rhetoric, academic AI research is increasingly driven by benchmark competition (Gururaja et al., 2023; Orr and Kang, 2024), often without questioning the assumptions embedded in said benchmarks (Raji et al., 2021). We encourage future researchers to examine what the universal aesthetics measurements used by companies and AI researchers actually measure.
5.2.2. Combining an Audit and Trace Ethnography
@@
In this work, we were able to examine upstream design decisions contributing to downstream biases by combining an audit and trace ethnography. Auditing is a core FAccT methodology, with the regex “audit*” appearing in nearly half of all FAccT publications. With some notable exceptions (Miceli et al., 2021; Marda and Narayan, 2020; Riley, 2024; Scheuerman, 2024; Passi and Barocas, 2019; Baack, 2024a), ethnography is far less common at FAccT. As Seaver (2017) argues, audits can help identify the existence of algorithmic biases while ethnographies can help explain where these biases may have originated by focusing on the people behind the algorithm. In addition to focusing on the harms of AI models themselves, responsible AI researchers are increasingly advocating for “studying up” the powerful people behind AI development and deployment (Barabas et al., 2020; Kawakami et al., 2024). In service of these goals, we advocate for combining audits and ethnographies as a socio-technical method for evaluating algorithmic systems.
@@
Trace ethnographies can be a powerful tool for studying AI development. While FAccT researchers do not need permission from developers to audit an algorithmic systems, one would likely need permission to conduct participant observations (Scheuerman, 2024). For example, prior ethnographic FAccT research has relied on access to field sites like tech companies (Scheuerman, 2024; Passi and Barocas, 2019) or a police headquarters (Marda and Narayan, 2020). However, maintaining access can limit critical scholarship (Scheuerman, 2024). Even in the case of “open” models, development practices may be opaque (Widder et al., 2024). We were only able to understand out how LAP was made by finding an explainer video in an old Discord message. Herein lies a lesson for future ethnographic FAccT research. As Seaver (2017) suggests, ethnographies of algorithms may require “scavenging” for diverse materials, including formal documentation, datasets, academic papers, and social media conversations. Nevertheless, this scavenging may pose ethical challenges. For example, studying an unlisted YouTube video—even if posted in a public Discord—could raise ethical concerns. We have chosen to include this video in our analysis due to LAION’s advocacy for data scraping and refusal to remove artists’ work from their datasets (Filikhina and Elsaß, 2025). That said, the “public-ness” of data does not always make its analysis ethical (Zimmer, 2010).
Finally, we caution that trace ethnographies are no substitute for more traditional ethnographic approaches, like extended participant observation and interviews. In our study, for instance, questions remain as to why the founder of LAION chose the datasets from which he built LAP. Likewise, it is not clear how members of LAION make sense of the widespread circulation of LAP. While our trace ethnography presents an account of how LAP was developed, extended participant observation would likely be needed to understand the culture of LAION within which LAP is situated. While we see potential for trace ethnographies to complement audits in FAccT research by embodying AI developers, this is not the only potential contribution of ethnographic analyses. Similar to Forsythe (2001)’s research on how expert system developers conceptualize “experts,” for example, we call for future ethnographic research into the how aesthetic evaluation researchers conceptualize “aesthetics.”
@@ 6. Conclusion
Generative AI models and algorithmic systems are increasingly mediating our media, art, and culture. Therefore, it is imperative to understand for whose taste these systems are designed. In this work, we audited the LAION-Aesthetics Predictor (LAP) model, finding that LAP reflects imperial, realist, and male aesthetic sensibilities. Through a trace ethnography of public materials, we found that the development of LAP reflects similar biases. In light of these findings, we call for greater attention to aesthetic biases as a source of representational harm and advocate for more pluralistic evaluations. Much like we examine the algorithmic gaze of LAP, we encourage future research into the cultural consequences of algorithmic “quality” assessment.
@@ Acknowledgements.
We thank Amy Bruckman, Rose Chang, Mary Gray, Anna Fang, Oliver Haimson, Shivani Kapania, Franchesca Spektor, Cella Sum, Sherry Tongshuang Wu and our reviewers for their feedback on this work.